### v1

@@ Abstract
Multimodal medical analysis combining image and tabular data has gained increasing attention. However, effective fusion remains challenging due to cross-modal discrepancies in feature dimensions and modality contributions, as well as the noise from high-dimensional tabular inputs. To address these problems, we present AMF-MedIT, an efficient Align-Modulation-Fusion framework for medical image and tabular data integration, particularly under data-scarce conditions. To harmonize dimension discrepancies and dynamically adjust modality contributions, we propose the Adaptive Modulation and Fusion (AMF) module, a novel modulation-based fusion paradigm with a streamlined architecture. We first derive the modulation objectives and introduce a modality confidence ratio, enabling the incorporation of prior knowledge into the fusion process. Then, the feature masks, density and leakage losses are proposed to achieve the modulation objectives. Additionally, we introduce FT-Mamba, a powerful tabular encoder leveraging a selective mechanism to handle noisy medical tabular data efficiently. Furthermore, interpretability studies are conducted to explore how different tabular encoders supervise the imaging modality during contrastive pretraining for the first time. Extensive experiments demonstrate that AMF-MedIT achieves a superior balance between multimodal performance and data efficiency while showing strong adaptability to incomplete tabular data. Interpretability analysis also highlights FT-Mamba’s capabilities in extracting distinct tabular features and guiding the image encoder toward more accurate and flexible attention patterns.
@@ 1 Introduction
Rrecently, multimodal learning that integrates medical images and tabular data has emerged as a critical research frontier in precision medicine. While computer-aided diagnosis based on images has been extensively investigated, the complementary role of tabular data in enhancing clinical decision-making is gaining attention. Such tabular data—typically derived from electronic health records (EHRs), laboratory tests, and standardized questionnaires—provide rich clinical context regarding the current symptoms and past medical history, which visual patterns alone cannot capture (Huang et al., 2020). Studies have consistently shown that integrating this clinical context improves image interpretation and supports more accurate diagnoses (Acosta et al., 2022).
@@
However, real-world applications of medical images and tabular data pose two challenges. First, significant disparities in feature dimension and confidence levels across modalities complicate the fusion process. (1) Dimension discrepancy: Image encoders require high-dimensional representations to retain anatomical details, whereas tabular encoders generate more compact features. For example, a ResNet-50 encoder extracts 2,048-dimensional global features, whereas a typical feature vector dimension of TabNet is 128 (Arik & Pfister, 2021). This imbalance in feature dimensions can lead to uneven parameter allocation during fusion, where higher-dimensional features tend to dominate in parameter space. (2) Confidence-level discrepancy: The reliability of each modality in multimodal fusion is influenced by the downstream task and data quality. When images provide definitive pathological signatures but tabular data suffer from missing values, an optimal model is expected to prioritize image evidence. Conversely, in predicting Ischemic Heart Disease (IHD), laboratory and lifestyle factors in tabular data often serve as more reliable prognostic indicators, with images playing a supplementary role (Zambrano Chaves et al., 2023). This disparity in confidence necessitates fusion strategies with adaptive capabilities to appropriately balance multimodal contributions. Second, medical tabular data inherently contain substantial noise, which hinders robust and accurate fusion (Borisov et al., 2024). These data consist of numerous features, many of which can be irrelevant to the target disease. For instance, among the selected 117 data fields for cardiac prediction in (Hager et al., 2023), over 60% have importance scores below 0.005 as measured by integrated gradients. Additionally, the prevalent problem of missing data further amplifies the noise of medical tabular data (Miao et al., 2023).
@@
These challenges, coupled with the persistent data scarcity in the medical domain, place greater demands on both fusion strategies and tabular representation learning, requiring a delicate balance between model capacity and data efficiency. Although the recent adaptation of Self-Supervised Learning (SSL) to the medical field has alleviated data scarcity in encoder training, core challenges in medical image-tabular fusion and tabular feature extraction remain unresolved. Naive concatenation fusion ignores both dimension and confidence disparities, whereas complex cross-attention mechanisms are data-hungry and computationally inefficient. In parallel, while architectures like Mamba (Gu & Dao, 2023) have advanced long-sequence modeling in natural language processing, medical tabular encoding still relies on either simplistic Multi-Layer Perceptrons (MLPs) or transformer-based models that require substantial training data to perform effectively.
@@
In this work, we aim to propose an efficient fusion paradigm that resolves cross-modal disparities through a streamlined architecture, while enhancing the tabular encoder under limited data conditions. Our method is based on three insights. First, inspired by contrastive learning, constructing a unified cross-modal representation space can mitigate representation discrepancies in a straightforward manner. However, unlike matching tasks, medical inference requires preserving modality-specific distinctions (Chen & Hong, 2024). This prompted us to construct a unified yet distinct cross-modal latent space by modulating only the dimension and scale of unimodal features while retaining essential information from the other modality. Second, given the scarcity of medical multimodal data, leveraging prior knowledge from preliminary experiments enables more efficient and interpretable fusion than directly learning cross-modal interactions from limited data. Therefore, we incorporate modality confidence priors into fusion to guide adaptive contribution allocation. Third, medical tabular data can be viewed as long feature sequences containing both relevant and irrelevant elements. Mamba’s selective and sequential modeling capabilities make it well-suited for processing such complex inputs (Gu & Dao, 2023). This motivates us to explore Mamba for medical tabular representation.
Based on the aforementioned insights, we present AMF-MedIT, an efficient Align-Modulation-Fusion framework for integrating medical image and tabular data. In the Align stage, we leverage contrastive learning to train unimodal encoders without annotations. In the Modulation stage, we present a novel paradigm to harmonize cross-modal discrepancies, generating fusion-compatible unimodal representations. Finally, in the Fusion stage, the modulated representations are integrated into a final model for downstream tasks.
Specifically, we propose an Adaptive Modulation and Fusion (AMF) module for unimodal feature modulation. To construct a unified yet distinct latent space for image and tabular features, we first derive the modulation objectives of similar feature densities and adaptive feature dimensions. Meanwhile, we define a modality confidence ratio $R_{conf}$ to incorporate prior information and guide adaptive fusion. In the concrete implementation, we propose feature masks to truncate features and density and leakage losses to compress information, thereby flexibly adjusting the feature dimensions within fixed network structures. Additionally, we introduce FT-Mamba, an adaptation of Mamba that enhances tabular feature extraction. Leveraging the selective State-Space Mechanisms (SSMs), FT-Mamba effectively captures discriminative features from noisy medical records without the need for large-scale training data. Furthermore, we conduct interpretability analysis to observe tabular encoders’ ability to preserve critical indicators and their supervisory role in imaging modality during contrastive learning. Substantial comparative experiments and interpretability studies further demonstrate the effectiveness of our method. The main contributions are summarized as follows.
  • We present AMF-MedIT, an efficient align-modulation-fusion framework for medical image and tabular data. It is designed to address key challenges in the medical domain, particularly in small-sample settings, and features a streamlined modulation-based fusion module coupled with a powerful tabular encoder.
@@
  • We propose the AMF module, a novel modulation-based fusion paradigm that mitigates dimension disparities and adaptively balances modality contributions by incorporating prior knowledge. We introduce FT-Mamba, to the best of our knowledge the first adaptation of Mamba for medical tabular data, aiming to enhance the feature extraction from noisy tabular data.
  • To the best of our knowledge, we conducted the first interpretability study on medical image-tabular contrastive learning, revealing how different tabular encoders impose distinct supervisory patterns on the imaging modality.
  • Experiments on both complete and incomplete tabular data demonstrate the superiority, computational efficiency, and adaptability of our method compared to mainstream fusion methods.
@@ 2.1 Self-Supervised Multimodal Learning
Self-supervised learning is proposed to extract meaningful representations from unlabeled data. In multimodal learning, training models still often require expensive human annotation (Krishnan et al., 2022). To mitigate this challenge, SSL has been applied to multimodal settings with various pretext tasks. CLIP utilizes image-text contrastive learning on massive web data, achieving remarkable zero-shot performance (Radford et al., 2021). Subsequent works also combine tasks such as Masked Language Modeling (MLM) and Image-Text Matching (ITM) to enhance cross-modal understanding (Li et al., 2021). Recently, SSL has been explored for image-tabular fusion (Hager et al., 2023; Du et al., 2024; Huang, 2023; Chen & Hong, 2024), but existing studies have not addressed specific challenges in the medical domain. Building on contrastive learning, our approach explicitly tackles the challenges in a medical context.
@@ 2.2 Multimodal Medical Image-Tabular Learning
Multimodal image-tabular learning exploits tabular data to facilitate visual task learning, which is widely used in the medical field (Huang et al., 2020). Several studies have employed late fusion, aggregating predictions from separate unimodal models for final decisions (Zambrano Chaves et al., 2023; Zhang et al., 2021). To facilitate deeper and earlier interactions between modalities, joint fusion has become more prevalent. Typically, features from different modalities are concatenated and then processed using fully connected networks (Lee et al., 2023; Vale-Silva & Rohr, 2021), transformer-based (Zheng et al., 2024, 2022), or traditional machine learning models (Zheng et al., 2020) for fusion. To enhance modality interaction, some studies employ cross-attention modules (Jia et al., 2024) to joint unimodal features. However, concatenation-based methods cannot reconcile the discrepancies between unimodal features, resulting in suboptimal fusion performance. Meanwhile, cross-attention methods rely entirely on full pairwise attention without leveraging prior knowledge, leading to high computational overhead, and making it difficult to achieve effective learning on small-scale medical datasets.
While SSL has advanced image-tabular learning, fusion strategies remain unchanged. Most studies still rely on feature concatenation (Hager et al., 2023; Huang, 2023; Painchaud et al., 2024) or cross-attention mechanisms (Zhou et al., 2023; Du et al., 2024; Chen & Hong, 2024; Chen et al., 2024). To address this limitation, we propose a novel fusion paradigm that flexibly regulates unimodal features within a streamlined structure, optimizing the trade-off between fusion performance and data efficiency.
@@ 2.3 Deep Learning for Tabular Data
Tabular data present unique challenges for deep learning (DL) models, requiring them to handle various data-related issues such as noise, imprecision, missing values, diverse attribute types, and varying value ranges (Borisov et al., 2024). Existing research on deep learning for tabular data provides diverse novel architectures. Gorishniy et al. (Gorishniy et al., 2021) compared the main deep architectures, including MLP, ResNet, and FT-Transformer, finding ResNet to be a simple yet effective baseline, whereas FT-Transformer serves as a universal architecture. Recently, Mamba has gained attention for leveraging selective SSMs to achieve linear-time complexity and memory efficiency (Gu & Dao, 2023). Thielmann et al. explored its application in the tabular domain, demonstrating Mamba’s feasibility for tabular data (Thielmann et al., 2024).
Despite this progress, tabular data processing in the medical domain remains relatively underdeveloped. Some studies use raw tabular data as input (Zheng et al., 2020; Pölsterl et al., 2021), while others employ MLP-like architectures for simple processing (Lee et al., 2023; Zheng et al., 2022; Vale-Silva & Rohr, 2021; Hager et al., 2023). However, such simplistic models often fall short of capturing meaningful representations from noisy and high-dimensional medical tabular data. Recent works have adopted transformer-based models to enhance tabular feature extraction (Zheng et al., 2024; Zhou et al., 2023; Du et al., 2024; Painchaud et al., 2024; Chen & Hong, 2024; Chen et al., 2024). Nevertheless, these models require large datasets for effective training and often encounter optimization difficulties on limited medical data. Inspired by FT-Transformer and Mambular, we propose a streamlined adaptation of the Mamba architecture for tabular data, specifically designed to handle the noisy nature of medical tabular data.
@@ 3.1 Overall Framework
Figure 1: The overall framework of AMF-MedIT. Modules outlined with black dashed borders are being trained in each stage. We employ FT-Mamba, a powerful and efficient encoder, for medical tabular data. Additionally, the AMF module is designed to modulate unimodal features, thereby facilitating effective fusion. Notably, $R_{conf}$ can be flexibly adjusted based on downstream requirements.
The overall framework of AMF-MedIT is shown in Fig. 1. The training pipeline comprises two stages: (a) Imaging-tabular pretraining and (b) Fine-tuning for adaptive modulation and fusion. In each stage, the image-tabular pairs, ($\boldsymbol{x}_{img}\in\mathbb{R}^{H\times W\times D}$, $\boldsymbol{x}_{tab}\in\mathbb{R}^{N}$) are fed into the corresponding unimodal encoders, $\boldsymbol{E}_{img}$ and $\boldsymbol{E}_{tab}$. Here, $H$, $W$, and $D$ denote the height, width, and number of channels of the input image, respectively, and $N$ represents the number of tabular features. We use ResNet-50 (He et al., 2016) as the image encoder and FT-Mamba as the tabular encoder, which will be introduced in Section 3.2. The encoders then generate the original unimodal embeddings, $\boldsymbol{z}_{img}\in\mathbb{R}^{D_{img}}$ and $\boldsymbol{z}_{tab}\in\mathbb{R}^{D_{tab}}$, where $D_{img}$ and $D_{tab}$ denote the dimensions of unimodal features.
@@
In the fine-tuning stage, we propose a simple AMF module for multimodal fusion. Guided by the proposed density and leakage losses, this module transforms the original unimodal features, $\boldsymbol{z}_{img}$ and $\boldsymbol{z}_{tab}$, into modulated features, $\boldsymbol{\tilde{z}}_{img}$ and $\boldsymbol{\tilde{z}}_{tab}$, which are then combined to form a cohesive fused feature, $\boldsymbol{z}_{fused}$. Finally, $\boldsymbol{z}_{fused}$ is fed to a Fully-Connected (FC) layer to generate the classification probability.

### v2

@@ Abstract
Multimodal medical analysis combining image and tabular data has gained increasing attention. However, effective fusion remains challenging due to cross-modal discrepancies in feature dimensions and modality contributions, as well as the noise from high-dimensional tabular inputs. To address these problems, we present AMF-MedIT, an efficient Align-Modulation-Fusion framework for medical image and tabular data integration, particularly under data-scarce conditions. Built upon a self-supervised learning strategy, we introduce the Adaptive Modulation and Fusion (AMF) module, a novel, streamlined fusion paradigm that harmonizes dimension discrepancies and dynamically balances modality contributions. It integrates prior knowledge to guide the allocation of modality contributions in the fusion and employs feature masks together with magnitude and leakage losses to adjust the dimensionality and magnitude of unimodal features. Additionally, we develop FT-Mamba, a powerful tabular encoder leveraging a selective mechanism to handle noisy medical tabular data efficiently. Extensive experiments, including simulations of clinical noise, demonstrate that AMF-MedIT achieves superior accuracy, robustness, and data efficiency across multimodal classification tasks. Interpretability analyses further reveal how FT-Mamba shapes multimodal pretraining and enhances the image encoder’s attention, highlighting the practical value of our framework for reliable and efficient clinical artificial intelligence applications.
@@ 1 Introduction
Rrecently, multimodal learning that integrates medical images and tabular data has emerged as a critical research frontier in precision medicine. While computer-aided diagnosis based on images has been extensively investigated, the complementary role of tabular data in enhancing clinical decision-making is gaining attention. Such tabular data—typically derived from electronic health records (EHRs), laboratory tests, and standardized questionnaires—provide rich clinical context regarding the current symptoms and past medical history, which visual patterns alone cannot capture [1]. Studies have consistently shown that integrating this clinical context improves image interpretation and supports more accurate diagnoses [2].
@@
However, real-world applications of medical images and tabular data pose two challenges. First, significant disparities in feature dimension and confidence levels across modalities complicate the fusion process. (1) Dimension discrepancy: Image encoders require high-dimensional representations to retain anatomical details, whereas tabular encoders generate more compact features. For example, a ResNet-50 encoder extracts 2,048-dimensional global features, whereas a typical feature vector dimension of TabNet is 128 [3]. This imbalance in feature dimensions can lead to uneven parameter allocation during fusion, where higher-dimensional features tend to dominate in parameter space. (2) Confidence-level discrepancy: The reliability of each modality in multimodal fusion is influenced by the downstream task and data quality. When images provide definitive pathological signatures but tabular data suffer from missing values, an optimal model is expected to prioritize image evidence. Conversely, in predicting Ischemic Heart Disease (IHD), laboratory and lifestyle factors in tabular data often serve as more reliable prognostic indicators, with images playing a supplementary role [4]. This disparity in confidence necessitates fusion strategies with adaptive capabilities to appropriately balance multimodal contributions. Second, medical tabular data inherently contain substantial noise, which hinders robust and accurate fusion [5]. These data consist of numerous features, many of which can be irrelevant to the target disease. For instance, among the selected 117 data fields for cardiac prediction in [6], over 60% have importance scores below 0.005 as measured by integrated gradients. Additionally, the prevalent problem of missing data further amplifies the noise of medical tabular data [7].
@@
These challenges, coupled with the persistent data scarcity in the medical domain, place greater demands on both fusion strategies and tabular representation learning, requiring a delicate balance between model capacity and data efficiency. Although the recent adaptation of Self-Supervised Learning (SSL) to the medical field has alleviated data scarcity in encoder training, core challenges in medical image-tabular fusion and tabular feature extraction remain unresolved. Naive concatenation fusion ignores both dimension and confidence disparities, whereas complex cross-attention mechanisms are data-hungry and computationally inefficient. In parallel, while architectures like Mamba [8] have advanced long-sequence modeling in natural language processing, medical tabular encoding still relies on either simplistic Multi-Layer Perceptrons (MLPs) or transformer-based models that require substantial training data to perform effectively.
@@
In this work, we aim to propose an efficient fusion paradigm that resolves cross-modal disparities through a streamlined architecture, while enhancing the tabular encoder under limited data conditions. Our method is based on three insights. First, inspired by contrastive learning, constructing a unified cross-modal representation space can mitigate representation discrepancies in a straightforward manner. However, unlike matching tasks, medical inference requires preserving modality-specific distinctions [9]. in which only the dimensionality and magnitude of unimodal features are aligned, while allowing the contributions and essential information specific to each modality to remain distinct. Second, given the scarcity of medical multimodal data, leveraging prior knowledge from preliminary experiments enables more efficient and interpretable fusion than directly learning cross-modal interactions from limited data. Therefore, we incorporate modality confidence priors into fusion to guide adaptive contribution allocation. Third, medical tabular data can be viewed as long feature sequences containing both relevant and irrelevant elements. Mamba’s selective and sequential modeling capabilities make it well-suited for processing such complex inputs [8]. This motivates us to explore Mamba for medical tabular representation.
Building on these insights, we present AMF-MedIT, an efficient Align–Modulation–Fusion framework for integrating medical images and tabular data. The framework first aligns the two modalities to obtain preliminary unimodal representations, then employs feature modulation to generate fusion-compatible unimodal representations, and finally fuses the modulated representations for downstream tasks. At its core, AMF-MedIT leverages the Adaptive Modulation and Fusion (AMF) module to ensure effective fusion. The modulation aims to reduce heterogeneity among unimodal features while adaptively adjusting the contributions of each modality, enabling robust and accurate integration across diverse downstream scenarios. Specifically, we first define a modality confidence ratio $r_{\text{conf}}$ to incorporate prior information regarding modality confidence, and derive two modulation objectives. For implementation, we design a streamlined, easy-to-train, and adaptive module architecture. The modulation objectives are achieved through feature masks that truncate features, complemented by magnitude and leakage losses to guide training.
Additionally, for multimodal representation learning and alignment, we introduce FT-Mamba, an adaptation of Mamba tailored for tabular feature extraction and pretrained using contrastive learning. Leveraging the selective State-Space Mechanisms (SSMs), FT-Mamba effectively captures discriminative features from noisy medical records without relying on large-scale training data. Meanwhile, the self-supervised contrastive learning strategy enables initial multimodal alignment in an unlabeled manner, allowing FT-Mamba to provide effective supervision for the image encoder.
We conducted extensive comparative experiments on both clean and noisy datasets. Results demonstrate that AMF-MedIT achieves a favorable balance between data efficiency and multimodal performance, showing strong adaptability to noise in both image and tabular modalities. Furthermore, ablation studies and interpretability analyses highlight the respective advantages of the AMF module in fusion and FT-Mamba in tabular feature extraction.
The main contributions are summarized as follows.
@@
  • We develop AMF-MedIT, an efficient framework for integrating medical images and tabular data. At its core, the AMF module is a novel modulation-based paradigm that mitigates dimension disparities and adaptively balances modality contributions by leveraging prior knowledge.
  • We introduce FT-Mamba, an adaptation of Mamba for medical tabular data that enhances feature extraction from noisy records, while employing self-supervised contrastive learning to enable unsupervised pretraining of image–table encoders.
  • We conduct extensive experiments under both clean and clinically noisy conditions, showing that AMF-MedIT achieves superior accuracy, robustness, and data efficiency compared with recent methods.
  • We provide in-depth ablation and interpretability analyses, revealing how the choice of tabular encoder influences multimodal contrastive pretraining.
@@ 2.1 Self-Supervised Multimodal Learning
Self-supervised learning is proposed to extract meaningful representations from unlabeled data. In multimodal learning, training models still often require expensive human annotation [10]. To mitigate this challenge, SSL has been applied to multimodal settings with various pretext tasks. CLIP utilizes image-text contrastive learning on massive web data, achieving remarkable zero-shot performance [11]. Subsequent works also combine tasks such as Masked Language Modeling (MLM) and Image-Text Matching (ITM) to enhance cross-modal understanding [12]. Recently, SSL has been explored for image-tabular fusion [6, 13, 14, 9], but existing studies have not addressed specific challenges in the medical domain. Building on contrastive learning, our approach explicitly tackles the challenges in a medical context.
@@ 2.2 Multimodal Medical Image-Tabular Learning
Multimodal image-tabular learning exploits tabular data to facilitate visual task learning, which is widely used in the medical field [1]. Several studies have employed late fusion, aggregating predictions from separate unimodal models for final decisions [4, 15]. To facilitate deeper and earlier interactions between modalities, joint fusion has become more prevalent. Typically, features from different modalities are concatenated and then processed using fully connected networks [16, 17, 33], transformer-based [18, 19], or traditional machine learning models [20] for fusion. To enhance modality interaction, some studies employ cross-attention modules [21] to joint unimodal features. However, concatenation-based methods cannot reconcile the discrepancies between unimodal features, resulting in suboptimal fusion performance. Meanwhile, cross-attention methods rely entirely on full pairwise attention without leveraging prior knowledge, leading to high computational overhead, and making it difficult to achieve effective learning on small-scale medical datasets.
While SSL has advanced image-tabular learning, fusion strategies remain unchanged. Most studies still rely on feature concatenation [6, 14, 22] or cross-attention mechanisms [23, 13, 9, 24]. To address this limitation, we propose a novel fusion paradigm that flexibly regulates unimodal features within a streamlined structure, optimizing the trade-off between fusion performance and data efficiency.
@@ 2.3 Deep Learning for Tabular Data
Tabular data present unique challenges for deep learning (DL) models, requiring them to handle various data-related issues such as noise, imprecision, missing values, diverse attribute types, and varying value ranges [5]. Existing research on deep learning for tabular data provides diverse novel architectures. Gorishniy et al. [25] compared the main deep architectures, including MLP, ResNet, and FT-Transformer, finding ResNet to be a simple yet effective baseline, whereas FT-Transformer serves as a universal architecture. Recently, Mamba has gained attention for leveraging selective SSMs to achieve linear-time complexity and memory efficiency [8]. Thielmann et al. explored its application in the tabular domain, demonstrating Mamba’s feasibility for tabular data [26].
Despite this progress, tabular data processing in the medical domain remains relatively underdeveloped. Some studies use raw tabular data as input [20, 27], while others employ MLP-like architectures for simple processing [16, 19, 17, 6]. However, such simplistic models often fall short of capturing meaningful representations from noisy and high-dimensional medical tabular data. Recent works have adopted transformer-based models to enhance tabular feature extraction [18, 23, 13, 22, 9, 24]. Nevertheless, these models require large datasets for effective training and often encounter optimization difficulties on limited medical data. Inspired by FT-Transformer and Mambular, we propose a streamlined adaptation of the Mamba architecture for tabular data, specifically designed to handle the noisy nature of medical tabular data.
@@ 3.1 Overall Framework
Figure 1: The overall framework of AMF-MedIT. (a) During the contrastive pretraining stage, unimodal encoders are trained without annotations to align image and tabular representations. We introduce FT-Mamba, a powerful tabular encoder for high-dimensional, noisy medical features. (b) In the modulation and fusion stage, we propose the AMF module, a novel modulation-based paradigm that harmonizes modality feature disparities and adaptively balances their contributions via prior-guided confidence ratio $r_{\text{conf}}$, along with the magnitude and leakage loss.
The overall framework of AMF-MedIT is shown in Fig. 1. The training pipeline comprises two stages: (a) Imaging-tabular pretraining and (b) Fine-tuning for adaptive modulation and fusion. In each stage, the image-tabular pairs, ($\boldsymbol{x}_{img}\in\mathbb{R}^{H\times W\times D}$, $\boldsymbol{x}_{tab}\in\mathbb{R}^{N}$) are fed into the corresponding unimodal encoders, $\boldsymbol{E}_{img}$ and $\boldsymbol{E}_{tab}$. Here, $H$, $W$, and $D$ denote the height, width, and number of channels of the input image, respectively, and $N$ represents the number of tabular features. We use ResNet-50 [28] as the image encoder and FT-Mamba as the tabular encoder, which will be introduced in Section 3.2. The encoders then generate the original unimodal embeddings, $\boldsymbol{z}_{img}\in\mathbb{R}^{D_{img}}$ and $\boldsymbol{z}_{tab}\in\mathbb{R}^{D_{tab}}$, where $D_{img}$ and $D_{tab}$ denote the dimensions of unimodal features.
@@
In the fine-tuning stage, we propose a simple AMF module for multimodal fusion. Guided by the proposed magnitude and leakage losses, this module transforms the original unimodal features, $\boldsymbol{z}_{img}$ and $\boldsymbol{z}_{tab}$, into modulated features, $\boldsymbol{\tilde{z}}_{img}$ and $\boldsymbol{\tilde{z}}_{tab}$, which are then combined to form a cohesive fused feature, $\boldsymbol{z}_{fused}$. Finally, $\boldsymbol{z}_{fused}$ is fed to a Fully-Connected (FC) layer to generate the classification probability.